OpenStudy (anonymous):

what are the steps i need to follow to solve this equation: 8 + x over 3 = 4?

OpenStudy (austinl):

\(\displaystyle \frac{8+x}{3}=4\) Do you have any thoughts?

OpenStudy (anonymous):

nope thnx

OpenStudy (austinl):

What do you do to cancel a division?

OpenStudy (anonymous):

i am not sure

OpenStudy (anonymous):

multiply both sides by 3